OpenStudy (anonymous):

Express your answer in lowest terms. -2/9x(-3/14)=?

OpenStudy (anonymous):

@Wendy.Ivette11714

OpenStudy (anonymous):

what its -2/9 times (-3/14)=?

OpenStudy (anonymous):

- 2/9) * (- 3/14) = [(- 2) * (- 3)]/[9 * 14] = (1)/(3 * 7) = 1/21

OpenStudy (anonymous):

thank yall